# LD-Scene: LLM-Guided Diffusion for Controllable Generation of Adversarial Safety-Critical Driving Scenarios

### v1

@@ IV-B Evaluation Metrics
IV-C Experimental Setup
Baseline. We evaluate our proposed method against existing baseline approaches for generating adversarial driving scenarios. Specifically, we compare with: our re-implementation of AdvSim [10], which optimizes the acceleration of a predefined adversarial vehicle to induce a collision, with initial states generated by SimNet [55]; and Strive [9], for which we utilize its open-source implementation. Strive employs a learned traffic model and performs adversarial optimization in the latent space.
Implementation Details. Our LD-Scene is implemented using the PyTorch framework and trained on four GeForce RTX 4090 GPUs for six hours. The diffusion model is trained for 200 epochs using the Adam optimizer with a learning rate of $5\times 10^{-4}$. The number of diffusion steps is set to 20. During the inference stage, we generate multiple candidate future trajectories for each non-ego agent in a given scene, with the number of test samples set to 10. The final trajectory is selected based on the one that minimizes the guidance loss. Additionally, both the code generator and debugger used in generating guidance utilize the GPT-4o model.
To ensure a fair evaluation of the controllability of the generated safety-critical scenarios, we adopt a standardized strategy for selecting the adversarial vehicle for each model. Specifically, we assign the adversarial vehicle as the one closest to the ego vehicle in the initial state of the scenario and ensure that it satisfies the required feasibility conditions. This selection strategy follows the approach used in Strive [9]; however, unlike Strive, the adversarial vehicle remains fixed throughout the scenario and does not change dynamically.
TABLE I: Overall performance comparison of baseline models on the nuScenes dataset. We compare our approach against AdvSim [10] and Strive [9] for safety-critical traffic simulation with a rule-based planner.
Model Adversariality Realism Efficiency
Adv-Ego Coll (%) $\uparrow$ Adv Acc $\uparrow$ Adv Offroad (%) $\downarrow$ Other Offroad (%) $\downarrow$ Adv-Other Coll (%) $\downarrow$ Other-Ego Coll (%) $\downarrow$ Other-Other Coll (%) $\downarrow$ Other Acc $\downarrow$ Infer time (s) $\downarrow$
AdvSim 24.72 0.90 15.60 14.85 0.56 0.91 0.11 0.38 338.35
Strive 22.69 0.88 18.94 16.64 0.90 1.08 0.05 0.39 609.72
LD-Scene 40.75 1.36 12.52 17.95 4.93 2.17 0.66 0.34 229.40
TABLE II: Ablation study on guidance settings. This table presents the impact of different guidance settings on model performance. The study evaluates the effects of different guidance loss components on adversarial effectiveness and realism in safety-critical traffic scenarios.
@@ IV-C Experimental Setup
TABLE III: Performance comparison under different adversarial levels.
       Adversarial Level        Adv-Ego Coll (%)        Adv Acc_lon        Adv Acc_lat        Adv Offroad (%)        FDD (m)
       Weak        30.63        0.58        0.69        10.81        59.52
       Medium        40.75        0.78        0.83        12.52        63.63
       Strong        39.33        0.82        0.92        14.53        78.44
IV-D Overall Performance
The quantitative performance of our approach and the baselines is presented in Table I. Compared to the baselines, the proposed LD-Scene demonstrates significant advantages in adversarial metrics while maintaining a high level of realism. Specifically, our LD-Scene achieves an Adv-Ego Coll (%) of 40.75%, which is substantially higher than the baseline models: AdvSim (24.72%) and Strive (22.69%). This highlights the superior capability of our model in controllably generating safety-critical driving scenarios that involve adversarial vehicles colliding with the ego vehicle. In terms of realism, the LD-Scene shows a lower Adv Offroad (%) of 12.52% compared to AdvSim (15.60%) and Strive (18.94%), indicating our model can better generate adversarial scenarios that are realistic and feasible within the road constraints. The LD-Scene also maintains an overall good performance across other metrics in realism. For instance, the Other Offroad (%) is 17.95%, which is comparable to the baseline models. Additionally, the collision rates involving other vehicles (Adv-Other Coll %, Other-Ego Coll %, and Other-Other Coll %) are all within reasonable ranges, further confirming the realistic nature of the generated scenarios. Furthermore, our diffusion-based framework significantly improves generation efficiency. LD-Scene achieves an average inference time of 229.40 seconds, substantially faster than the test-time optimization methods employed by AdvSim (338.35 s) and Strive (609.72 s). These results collectively highlight the effectiveness of LD-Scene in generating diverse, realistic, and computationally efficient safety-critical traffic scenarios.
Overall, these results demonstrate the capability of our framework to generate realistic, diverse, safety-critical driving scenarios, which can serve as valuable stress tests for autonomous vehicle systems.
IV-E Ablation Study
The ablation study results presented in Table II highlight the effectiveness of each loss component to the overall generation performance. Specifically, with the integration of all three guidance losses (Other-real Guidance, Adv-real Guidance, and Adv Guidance), the model shows superior performance compared to a baseline diffusion model without any guidance. The Adv Guidance component is particularly critical for increasing the collision rate between an adversarial vehicle and the ego vehicle, achieving an Adv-Ego Coll (%) of 39.68%, far exceeding the baseline (1.26%). Additionally, the Adv-real Guidance significantly lowers the probability of adversarial vehicles going off-road, reducing the Adv Offroad (%) from 13.15% (without guidance) to 10.68%, thereby enhancing the realism and feasibility of adversarial trajectories. Meanwhile, the Other-real Guidance improves the feasibility of non-adversarial vehicle trajectories by decreasing the Other Offroad (%) to 17.95% and maintaining lower collision rates involving other vehicles, such as Adv-Other Coll (%) at 4.93% and Other-Other Coll (%) at 0.66%. Although this slightly increases the Adv Offroad (%), it contributes to an overall enhancement in realism. Consequently, the proposed LD-Scene achieves an optimal Adv-Ego Coll (%) of 40.75% while keeping the Adv Offroad (%) at a reasonable level of 12.52%, confirming its effectiveness in generating both highly adversarial and realistic safety-critical driving scenarios.
IV-F Controllability Study
IV-F1 Controllable Adversarial Level
@@
We further investigate the controllability of our LD-Scene in terms of generating safety-critical driving scenarios with varying levels of adversarial intensity. Three adversarial levels, weak, medium, and strong, are introduced for the study, with the associated user queries defined as follows:
@@
The performance comparison under different adversarial levels is presented in Table III, which demonstrates that our model can infer the desired adversarial intensity from the user query and generate corresponding safety-critical driving scenarios. At the weak adversarial level, the model achieves a notable Adv-Ego Coll (%) of 30.63%, indicating a good performance in generating low-speed collisions between the adversarial vehicle and the ego vehicle. The values of Adv Acc_lon and Adv Acc_lat suggest that the generated trajectories are relatively less aggressive, and the low Adv Offroad (%) of 10.81% manifests a high level of realism. However, the FDD (m) value of 59.52 indicates a limited diversity in the generated scenarios. At the medium level, the Adv-Ego Coll (%) increases to 40.75%, and FDD rises to 63.63, indicating enhanced adversarial behavior and broader scenario diversity. However, this comes with a slightly reduced realism, as indicated by the higher Adv Offroad (%) of 12.52%. At the strong adversarial level, the model achieves the highest FDD of 78.44, showing exceptional scenario diversity. Meanwhile, the Adv-Ego Coll (%) of 39.33% is comparable to the medium level (40.75%), showing that the strong level does not necessarily lead to a higher collision rate between the adversarial vehicle and ego vehicle across the NuScenes dataset. This could be because high-speed overtaking maneuvers may not always result in increased collision opportunities, especially in low-speed interaction scenarios where the timing for generating adversarial collisions can be easily missed.
@@
In summary, our LD-Scene effectively controls the adversarial level, allowing users to generate safety-critical driving scenarios tailored to their specific needs. From weak to strong adversarial levels, the model balances adversarial collision rates, scene diversity, and realism, providing a versatile tool for enhancing the robustness of autonomous driving systems.
IV-F2 Controllable Adversarial Behavior
@@
We conduct three case studies on different user queries for generating adversarial safety-critical scenarios, demonstrating our LD-Scene’s capability to synthesize diverse and controllable adversarial behaviors, as illustrated in Fig. 4. Each case encompasses two example scenarios presented below the dialog. In case (a), the user query specifies a normal collision, leading LD-Scene to generate a loss function that minimizes the relative distance between the ego and adversarial vehicles. In case (b), the user query further includes a high-speed requirement, prompting LD-Scene to introduce an additional loss function to encourage high velocity. The associated scenarios illustrate the adversarial vehicle accelerating to overtake the ego vehicle before colliding, effectively simulating reckless high-speed maneuvers. Regarding case (c), the user query requests a sharp turn leading to a collision, and LD-Scene responds by generating a corresponding turn-related loss function. In both the example scenarios, the adversarial vehicle abruptly swerves off its path, ultimately resulting in a collision. These results showcase that our LD-Scene can dynamically generate safety-critical scenarios based on different user queries, which provides a flexible and controllable framework for evaluating autonomous vehicle performance under diverse adversarial situations.
IV-G Analysis of the Impact of Key Parameters
@@
We analyze the impact of two key parameters, namely the number of samples and the number of diffusion steps, on adversarial scenario generation. The evaluation is based on three criteria: adversariality (Adv-Ego collision rate), realism (Adv Offroad collision rate), and efficiency (time consumption).
@@
V Conclusion
In this paper, we present LD-Scene, an LLM-guided diffusion framework for the controllable generation of adversarial safety-critical driving scenarios. By integrating LDMs with LLM-enhanced guidance, our approach enables flexible, user-friendly scenario generation while ensuring both realism and adversarial effectiveness. Our method leverages CoT reasoning for structured guidance loss generation and incorporates an automated debugging module to enhance the reliability and stability of the generated guidance. Extensive experiments conducted on the nuScenes dataset demonstrate that our LD-Scene outperforms existing adversarial scenario generation baselines in terms of adversarial effectiveness, realism, and diversity. Moreover, our framework allows natural language-based customization, making it accessible to users without extensive domain expertise. The controllability studies further demonstrate that our LD-Scene effectively modulates both the adversarial level and some specific adversarial behaviors, facilitating a more rigorous evaluation of AV performance under diverse safety-critical scenarios.

### v2

@@ 4.4 Overall Performance
4.5 Ablation Study
4.5.1 Effectiveness of the Guidance Components
Table 2 demonstrates the contribution of each guidance component to the overall generation performance. Specifically, with the integration of all three guidance losses (Other-real Guidance, Adv-real Guidance, and Adv Guidance), the model shows superior performance compared to a baseline diffusion model without any guidance. The Adv Guidance component is particularly critical for increasing the collision rate between an adversarial vehicle and the ego vehicle, achieving an Adv-Ego Coll (%) of 39.68%, far exceeding the baseline (1.26%). Additionally, the Adv-real Guidance lowers the probability of adversarial vehicles going off-road, reducing the Adv Offroad (%) from 13.15% to 10.68%, thereby enhancing the realism and feasibility of adversarial trajectories. Meanwhile, the Other-real Guidance improves the feasibility of non-adversarial vehicle trajectories by decreasing the Other Offroad (%) to 17.95% and maintaining lower collision rates involving other vehicles, such as Adv-Other Coll (%) at 4.93% and Other-Other Coll (%) at 0.66%. Consequently, the proposed LD-Scene achieves an optimal Adv-Ego Coll (%) of 40.75% while keeping the Adv Offroad (%) at a reasonable level of 12.52%, confirming its effectiveness in generating both highly adversarial and realistic safety-critical driving scenarios.
@@ 4.5.1 Effectiveness of the Guidance Components
Table 2: Ablation study on guidance settings. This table presents the impact of different guidance settings on model performance. The study evaluates the effects of different guidance loss components on adversarial effectiveness and realism in safety-critical traffic scenarios.
Figure 5: Ablation study on the effectiveness of the debugger module. (a): Success rates of guidance code execution with and without the debugger. The debugger significantly improves success rates across all LLMs. (b): Average token input and output statistics per model. (c): Total cost of each model.
4.5.2 Effectiveness of the Debugger Module
To validate the stability improvements afforded by the debugger module in our LD-Scene, we compare the execution success rates of the generated guidance functions for different LLM models under both with and without debugger settings. We evaluate a total of 500 user queries, which are automatically generated by GPT-4o. As shown in Fig. 5(a), the debugger significantly improves success rates for all evaluated LLMs. For example, GPT-4o’s success rate increases from 69.4% to 95.0%, achieving highly stable generation. These results confirm that the debugger module can effectively detect and correct generation errors, ensuring reliable guidance function generation. In addition, we assess token consumption and total cost for different LLM models in Fig. 5(b) and Fig. 5(c), respectively. These results confirm that the debugger module offers substantial reliability improvements at a reasonable computational and financial expense. These insights further provide practical guidance for researchers in selecting suitable LLM models for similar tasks.
4.6 Controllability Study
4.6.1 Controllable Adversarial Level
@@
We further investigate the controllability of our LD-Scene in terms of generating safety-critical driving scenarios with varying levels of adversarial intensity. Three adversarial levels, weak, medium, and strong, are introduced for the study, and the three user query examples used in this experiment are as follows:
@@
The performance comparison under different adversarial levels is presented in Table 3, which demonstrates that our model can infer the desired adversarial intensity from the user query and generate corresponding safety-critical driving scenarios. At the weak adversarial level, the metrics such as TTC, Adv Acc_lon, and Adv Acc_lat indicate less aggressive driving behaviors. In contrast, both the Medium and Strong levels exhibit progressively more aggressive maneuvers, as evidenced by shorter TTC and increased acceleration values, thereby validating the controllability of adversarial behavior intensity. As illustrated in Fig. 6(a), the TTC distribution shows progressively narrower peaks and lower medians from weak to strong levels. Meanwhile, Fig. 6(b) and Fig. 6(c) present the longitudinal and lateral acceleration distributions, respectively, which exhibit increasingly broader spreads toward higher acceleration values. These visualizations more clearly demonstrate the model’s ability to modulate scenario aggressiveness across multiple safety-critical dimensions. An additional observation is that the Adv-Ego Coll (%) at the Strong level(39.33%) is comparable to that at the medium level (40.75%), showing that the strong level does not necessarily lead to a higher collision rate between the adversarial vehicle and ego vehicle across the NuScenes dataset. This could be because high-speed overtaking maneuvers may not always result in increased collision opportunities, especially in low-speed interaction scenarios where the timing for generating adversarial collisions can be easily missed.
@@
In summary, our LD-Scene provides effective control over the adversarial level, allowing users to generate safety-critical driving scenarios tailored to their specific needs. From weak to strong adversarial levels, the model serves as a practical tool for testing the performance and safety of autonomous driving systems under diverse risk conditions.
Adversarial Level Adv-Ego Coll (%) TTC (s) Adv Acc_lon (m/s2) Adv Acc_lat (m/s2) Adv Offroad (%)
Weak 30.63 2.06 1.02 0.69 10.81
Medium 40.75 1.98 1.09 0.83 12.52
Strong 39.33 1.91 1.20 0.92 14.53
Table 3: Performance comparison under different adversarial levels.
Figure 6: Visualization of results under different adversarial levels.
4.6.2 Controllable Adversarial Behavior
@@
We conduct three case studies on different user queries for generating adversarial safety-critical scenarios, demonstrating our LD-Scene’s capability to synthesize diverse and controllable adversarial behaviors, as illustrated in Fig. 7. Each case encompasses two example scenarios presented below the dialog. In case (a), the user query specifies a normal collision, leading LD-Scene to generate a loss function that minimizes the relative distance between the ego and adversarial vehicles. In case (b), the user query further includes a high-speed requirement, prompting LD-Scene to introduce an additional loss function to encourage high velocity. The associated scenarios illustrate the adversarial vehicle accelerating to overtake the ego vehicle before colliding, effectively simulating reckless high-speed maneuvers. Regarding case (c), the user query requests a sharp turn leading to a collision, and LD-Scene responds by generating a corresponding turn-related loss function. In both example scenarios, the adversarial vehicle abruptly swerves off its path, ultimately resulting in a collision. These results showcase that our LD-Scene can dynamically generate safety-critical scenarios based on different user queries, which provides a flexible and controllable framework for evaluating autonomous vehicle performance under diverse adversarial situations.
4.7 Analysis of the Impact of Key Parameters
@@
We analyze the impact of two key parameters, namely the number of samples and the number of diffusion steps, on safety-critical scenario generation. The evaluation is based on three criteria: adversariality (Adv-Ego collision rate), realism (Adv Offroad collision rate), and efficiency (time consumption).
@@
5 Conclusion
In this paper, we present LD-Scene, an LLM-guided diffusion framework for the controllable generation of adversarial safety-critical driving scenarios. By integrating LDMs with LLM-enhanced guidance, our approach enables flexible, user-friendly scenario generation while ensuring both realism and adversarial effectiveness. Our method leverages CoT reasoning for structured guidance loss generation and incorporates an automated debugging module to enhance the reliability and stability of the generated guidance. Extensive experiments conducted on the nuScenes dataset demonstrate that our LD-Scene outperforms existing adversarial scenario generation baselines in terms of adversarial effectiveness, realism, and efficiency. Moreover, our framework allows natural language-based customization, making it accessible to users without extensive domain expertise. The controllability studies further demonstrate that our LD-Scene effectively modulates both the adversarial level and some specific adversarial behaviors, facilitating a more rigorous evaluation of AV performance under diverse safety-critical scenarios.